# Out-of-equlibrium inference of stochastic model parameters through population data from generic consumer-resource dynamics

### v1

@@ II.2 Out-of-equilibrium probability distribution
where $p_{A}=\nu/(\nu+Q)$ and $p_{j}=\theta_{j}/(\nu+Q)$, which are precisely the probabilities given by Eqs. (32) and (33) in the particular case of equal handling times. Therefore, in this limit we obtain the generating function (61) of the multinomial distribution $\mathcal{M}(n_{A}^{0},\{p_{A},p_{1},\dots,p_{S}\})$ with $n_{A}^{0}$ trials. In sum, what we have shown here is that, in the case of equal handling times, and starting by an initial population formed solely by $n_{A}^{0}$ searching consumers, the joint probability distribution of all behavioral types is multinomial in the transient as well as at stationarity. Consistently, the probabilities $p_{j}(t)$ derived in the out-of-equilibrium regime, given by Eq. (62), satisfy $p_{j}(t)\to p_{j}$ ($j=1,\dots,S$) and $p_{0}(t)\to p_{A}$ in the limit $t\to\infty$, for $p_{A}$ and $p_{j}$ given by Eqs. (32) and (33), respectively.
III Model identifiability
Now we focus on the problem of model identifiability. Is it possible to identify all the parameters that define chemical reactions based only on feeding experiments, in which one may have access to the population abundances of each behavioral type? We distinguish between two situations: (i) abundance data is available for different independent replicas of the system at equilibrium, and (ii) abundances for different behavioral types are available for out-of-equilibrium independent replicas of the system, taken at different finite times. Based on the previous analytical results, in this section we show that reaction parameters are not fully identifiable in the first situation (one can recover parameter ratios), but they are fully identifiable in the out-of-equilibrium scenario. We will illustrate this issue through in silico experiments, so we can properly apply the mathematical results derived above.
@@ III Model identifiability
First, consider an in silico feeding experiment with $n_{A}^{0}$ individual consumers feeding on $S$ different resources at controlled resource densities ($x_{R_{i}}$ constant, $i=1,\dots,S$), under Holling type II stochastic dynamics —cf. reaction schemes in Eqs. (2)–(3). The parameter values that we would like to estimate from abundance data are: attack rate $\alpha_{i}$ for each resource, as well as the inverse of average handling time over each resource, $\nu_{i}$ (relaxation rate). Assume a fixed system volume $N$. Suppose that the system is sampled at a time large enough for it to have reached stationarity. Suppose we have a detection device that can record the number of free consumers, $n_{A}$, as well as the number of consumers handling on each resource $i$, $n_{i}$, for $i=1,\dots,S$. We replicate the experiment a number of times $T$ to obtain independent stochastic realizations of the system’s steady state. Our observational data is the following $S\times T$ matrix listing the observations of the number of predators handling on each resource,
$M=\begin{pmatrix}n_{1,1}&n_{1,2}&\cdots&n_{1,T}\\ \vdots&\vdots&\ddots&\vdots\\ n_{S,1}&n_{S,2}&\cdots&n_{S,T}\end{pmatrix},$ (67)
which gathers $T$ independent observations or stochastic realizations. Free consumer abundances in each replica are determined by the relation $n_{A,k}=n_{A}^{0}-\sum_{j=1}^{S}n_{j,k}$.
As explained, Eq. (26) allows to calculate the joint (multinomial) probability distribution at stationarity for $n_{A}^{0}$ trials and probabilities given by Eqs. (32) and (33), $P_{\bm{n}}(\bm{\Theta})$. This function gives the probability of observing a population vector $(n_{A},\bm{n})$, where $n_{A}=n_{A}^{0}-\sum_{j=1}^{S}n_{j}$, given a model parameter vector, $\bm{\Theta}:=(\alpha_{1},\dots,\alpha_{S},\nu_{1},\dots,\nu_{S})$ —notice that resource densities, $x_{R_{i}}=n_{R_{i}}/N$, the system volume, $N$, and the total number of consumers, $n_{A}^{0}$, are kept constant and therefore are not considered as parameters in $\bm{\Theta}$. Given certain parameter values, assuming that replicates are independent realizations of the stochastic chemical reaction scheme, we can assign a multiplicative likelihood to the observed matrix, $M$:
$\mathcal{L}(\bm{\Theta}|M)=\prod_{i=1}^{T}P_{\bm{n}_{i}}(\bm{\Theta}),$ (68)
where $\bm{n}_{i}:=(n_{1,i},\dots,n_{S,i})$ is the observed vector of handling predator abundances (corresponding to the $i$-th column of matrix $M$). The log-likelihood function,
$\log\mathcal{L}(\bm{\Theta}|M)=\sum_{i=1}^{T}\log P_{\bm{n}_{i}}(\bm{\Theta}),$ (69)
@@
Figure 2: Parameter estimation at equilibrium. We generated a single realization of the data matrix $M$, including $T=100$ observations, using Gillespie’s stochastic simulation algorithm for $\alpha=5$, $\nu=10$, and $n_{A}^{0}=100$ individual consumers feeding on $S=5$ resources with densities $\bm{x_{R}}=(0.9,0.7,0.5,0.3,0.1)$. Left: for that synthetic dataset, we run a L-BFGS-B optimization algorithm for the log-likelihood function (69) starting from different initial conditions, and we recorded the values of $(\alpha^{\star},\nu^{\star})$ at which the maximum was reached. Optimal parameter values follow a perfect straight line. The distributions of $\alpha^{\star}$ and $\nu^{\star}$ are compatible with uniform distributions, and the optimal values of $\log\mathcal{L}$ keep constant up to round-off errors. Overall, this indicates a line of infinite optima. Right: we replicated the previous experiment 300 times and, for each replica, we estimated the ratio $\nu/\alpha$ by fitting the slope of the straight line. We obtained the distribution in the right panel, which is compatible with a Gaussian with mean $\mu=2$ and variance extracted from data (Kolmogorov-Smirnov test, $p=0.96$). The vertical, dashed line indicates the true ratio.
The steady-state, joint probability distribution (26) depends only on the dimensionless rate quotients $\theta_{i}/\nu_{i}=x_{R_{i}}\alpha_{i}/\nu_{i}$, for $i=1,\dots,S$. Because $x_{R_{i}}$ are kept constant under the assumption of chemostatic conditions, the relevant, dimensionless chemical reaction rates are $\widetilde{\nu}_{i}:=\nu_{i}/\alpha_{i}$, i.e., the relaxation rate of the compound $X_{AR_{i}}$ in units of the predator’s attack rate on resource $i$. In sum, the likelihood function of the multi-species Holling type II case depends on reaction rates only through dimensionless ratios $\widetilde{\nu}_{i}$, $i=1,\dots,S$. These dimensionless rates are the ones that can be truly estimated from data at stationarity. The true reaction rates $\alpha_{i}$ and $\nu_{i}$ are not identifiable if the in silico experiment is conducted at equilibrium.
@@
Figure 2 shows why the model is not identifiable at equilibrium. In silico data (i.e., abundances of each behavioral type) were generated by stochastic simulation of a system with $S=5$ resources (with fixed densities $x_{R_{i}}$) consumed by $n_{A}^{0}$ predator individuals, all of them having equal attack rates ($\alpha:=\alpha_{1}=\dots=\alpha_{S}$) and equal relaxation rates ($\nu:=\nu_{1}=\dots=\nu_{S}$). A number $T$ of independent stochastic simulations, generated with the Gillespie algorithm [32], started with the initial conditions $n_{A}=n_{A}^{0}$ and $n_{i}=0$ ($i=1,\dots,S$) at $t=0$, each of which were stopped up to time $t_{\text{end}}=50\tau$, enough to reach stationarity, $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ being the characteristic time of the dynamics —see Eq. (62). Therefore, at stationarity we recorded $T$ abundance vectors $\bm{n}$ to form matrix $M$ column-wise. We maximized the log-likelihood function using the L-BFGS-B optimization scheme [33], which was used for numerical optimization throughout the contribution. Each realization of the data matrix $M$ leads to a line of maximum likelihood in the $(\alpha,\nu)$ space (see also Fig. 3, left panel), whose slope can vary over different replicas but can be determined precisely, leading to a consistent, unbiased estimation of the ratio $\widetilde{\nu}=\nu/\alpha$ (Fig. 2, right panel). At equilibrium, only parameter ratios are identifiable from data.
@@
From Eqs. (34)–(35), it is immediate to check that
$\widetilde{\nu}=\frac{\nu}{\alpha}=\frac{\langle n_{A}\rangle}{\langle n_{AR}\rangle}\sum_{i=1}^{S}x_{R_{i}},$ (70)
@@
with $\langle n_{AR}\rangle:=\left\langle\sum_{i=1}^{S}n_{i}\right\rangle=\sum_{i=1}^{S}\langle n_{i}\rangle$ being the average total number of handling consumers. In Appendix B we show that the likelihood function does exhibit a line of infinite critical points with slope given by (70), in which quantities $\langle n_{A}\rangle$ and $\langle n_{AR}\rangle$ are substituted by averages across temporal samples, $\langle n_{A}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{A,k}$ and $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}\sum_{i=1}^{S}n_{i,k}$, respectively. Therefore, our numerical experiment (Fig. 2, left) is consistent with the expected optimal behavior of the log-likelihood.
@@
Figure 3: Heatmaps of negative log-likelihood. Left: representation of $-\log\mathcal{L}$, cf. Eq. (68), in the $(\alpha,\nu)$ space, obtained for a single model realization when samples are taken at equilibrium. A line of infinite minima is obtained. Right: a single out-of-equilibrium model replica yields a heatmap for $-\log\mathcal{L}$, according to Eq. (72). Now, $T=100$ sampling times were uniformly distributed over the interval $(0,10\tau)$, $\tau$ being the characteristic time of the dynamics. The arrow marks the true pair $(\alpha,\nu)=(5,10)$ used for data generation, and the white point marks the minimum of $-\log\mathcal{L}$. In both panels, parameter values are the same as in Fig. 2.
If the experiment is done out of equilibrium, the problem of model identifiability is solved. Consider a second in silico experiment in which samples of the system state are taken before reaching the steady state. Again, in the experiment there are $n_{A}^{0}$ individual consumers feeding on $S$ resources, under chemostatic conditions. Assume that the system is driven by a Holling type II stochastic dynamics. The system is sampled at finite times $t_{i}$, before reaching stationarity, times at which population abundances $n_{A}(t_{i})$ and $n_{j}(t_{i})$, for $j=1,\dots,S$, are recorded. Since the system is out-of-equilibrium and we are evaluating transient abundances, it is important to specify the initial condition of each realization: we assume that every replica of the experiment starts initially with $n_{A}(0)=n_{A}^{0}$ free consumers, and no consumer is handling any prey. To illustrate how the model is fully identifiable out of equilibrium, we make the hypothesis that $\nu_{1}=\dots=\nu_{S}$, i.e., we consider from now on that consumer average handling times are equal across resources. Thus, we can use our previous derivation of the multinomial distribution for an arbitrary time: the probability of a particular system configuration is given by Eq. (26) with probabilities $p_{j}(t)$ given by Eq. (62).
We replicate the experiment a number of times $T$, each replica simulating the stochastic dynamics (using Gillelpie’s method, [32]) with $n_{A}^{0}$ free predators initially, and record abundances at time $t_{i}$ for each realization $i=1,\dots,T$. As shown in Appendix B, it is important that at least two sampling times are different (i.e., $t_{j}\neq t_{k}$ for some $j\neq k$). Otherwise the model is not identifiable. In practice, from an empirical point of view, it is very unlikely that independent system samples are obtained at precisely equal times.
Thus, we have $T$ independent stochastic realizations of the system state, each replica is an observation at different times before reaching the steady state. Now, the observational data is the following matrix
$M(\{t_{i}\})=\begin{pmatrix}n_{1,1}(t_{1})&n_{1,2}(t_{2})&\cdots&n_{1,T}(t_{T})\\ \vdots&\vdots&\ddots&\vdots\\ n_{S,1}(t_{1})&n_{S,2}(t_{2})&\cdots&n_{S,T}(t_{T})\end{pmatrix},$ (71)
where columns comprise out-of-equilibrium vectors $\bm{n}_{i}$ reporting observed abundances of handling predators at each sampling time. Free predators abundances are computed as $n_{A,i}(t_{i})=n_{A}^{0}-\sum_{j=1}^{S}n_{j,i}(t_{i})$. Again, the likelihood function is multiplicative, but we have to use the time dependent multinomial probabilities to assign a likelihood to the parameter vector $\bm{\Theta}$ given the data matrix $M(\{t_{i}\})$:
$\mathcal{L}(\bm{\Theta}|M(\{t_{i}\}))=\prod_{i=1}^{T}P_{\bm{n}_{i}}(\bm{\Theta},t_{i}),$ (72)
where $P_{\bm{n}_{i}}(\bm{\Theta},t_{i})$ is given by Eq. (26) and (62), and in this case the parameter vector is $\bm{\Theta}:=(\alpha_{1},\dots,\alpha_{S},\nu)$.
@@
Figure 4: Distributions of parameter estimates. (a) The estimation procedure was repeated three times, for $n_{A}^{0}\in\{20,100,500\}$, using $T=20$ samples in simulated data. Upper panels show the distributions of $\alpha^{\star}$ and $\nu^{\star}$, which tend to concentrate around true values (dashed lines) as $n_{A}^{0}$ increases. Parameter estimates show bias for low $n_{A}^{0}$ values, and the bias reduces as $n_{A}^{0}$ increase. This phenomenon is more clearly observed in the density plots represented in lower panels for $(\alpha^{\star},\nu^{\star})$, the lighter the color indicating the higher density of estimated pairs. (b) Same as (a) for $T=100$ temporal samples. Compared to panel (a), we observe that variability is decreased as $T$ increases. In both panels, remaining parameters are the same as in Fig. 2. $3000$ estimation experiments were conducted to obtain distributions.
Observe how probability parameters $p_{j}(t)$ in Eq. (62) depend on $\bm{\Theta}$. Now it is clear how rates $\alpha_{i}$ and $\nu$ can be properly inferred themselves from data, because they appear in $p_{j}(t)$ via $S$ ratios $\alpha_{i}/\nu$ and a sum $\nu+\sum_{k=1}^{S}x_{R_{k}}\alpha_{k}$. At the maximum of the likelihood function, the optimization method would yield numerical values for these $S+1$ parameter combinations, let’s say $\widetilde{\nu}_{i}:=\nu/\alpha_{i}$ and $\widetilde{\xi}:=\nu+\sum_{k=1}^{S}x_{R_{k}}\alpha_{k}$. They form a non-linear system of $S+1$ equations with $S+1$ unknowns, from which one can determine the maximum-likelihood (ML) estimates of the rates, $\alpha_{i}^{\star}$ and $\nu^{\star}$, as
$\nu^{\star}=\widetilde{\xi}\left(1+\sum_{k=1}^{S}\frac{x_{R_{k}}}{\widetilde{\nu}_{k}}\right)^{-1},\quad\alpha_{i}^{\star}=\frac{\nu^{\star}}{\widetilde{\nu}_{i}}.$ (73)
Therefore, the ML estimation method will produce proper estimates for rates $\alpha_{i}$ and $\nu$ themselves: the problem of model identifiability if solved if the experiment is conducted out of equilibrium and, importantly, at least two sampling times are different (see Appendix B).
@@
Figure 3 compares the log-likelihood functions at equilibrium, Eq. (68), and out of equilibrium, Eq. (72), when attack and relaxation rates are equal across resources, $\alpha=\alpha_{1}=\dots=\alpha_{S}$ and $\nu=\nu_{1}=\dots=\nu_{S}$. A parameter screening in the $(\alpha,\nu)$ plane shows a line of infinite maxima $\nu^{\star}\propto\alpha^{\star}$ at equilibrium, evidencing the results in Fig. 2 and Appendix B. This proves the lack of model identifiability at equilibrium. On the contrary, when samples are taken out of equilibrium, a single likelihood maximum value arises and parameters can be truly inferred from data. In the out-of-equilibrium case, $T=100$ temporal samples were taken, and sampling times $\{t_{i}\}$ were drawn uniformly over the interval $(0,10\tau)$, $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ being the characteristic dynamical time scale. In this case, a unique maximum is found for the log-likelihood function. The accuracy of the estimates $\alpha^{\star}$ and $\nu^{\star}$ depends on data quality. For example, if the number of samples $T$ taken is large, then it is expected that estimates will be close to actual values. Similarly, if the overall number of individuals $n_{A}^{0}$ monitored in the experiment is large, then better estimates for model parameters are expected.
@@
Usually, we only have access to one single realization of empirical observations. Then we can obtain ML parameter estimates as well as confidence intervals following the log-likelihood ratio test [34]. We illustrate this procedure using the same data matrix as in Fig. 3b, for which the ML parameter estimates are $(\alpha^{\star},\nu^{\star})=(5.24,10.65)$. We obtained log-likelihood profiles shown in Fig. 5 by minimizing negative log-likelihood over $\nu$ for fixed $\alpha$, $\ell_{1}(\alpha):=-\min_{\nu}\log\mathcal{L}(\alpha,\nu|M(\{t_{i}\}))$ (left panel), and minimizing over $\alpha$ for fixed $\nu$, $\ell_{2}(\nu):=-\min_{\alpha}\log\mathcal{L}(\alpha,\nu|M(\{t_{i}\}))$ (right panel). Then, the $95\%$ confidence interval for $\alpha$ can be defined as the set of values satisfying $|\ell_{1}(\alpha)-\ell_{1}(\alpha^{\star})|\leq\chi^{2}_{1,0.95}/2$, $\chi^{2}_{1,0.95}$ being the $95\%$ percentile of the chi-squared distribution with one degree of freedom. Similarly, the $95\%$ confidence interval for $\nu$ is defined as the set of values satisfying $|\ell_{2}(\nu)-\ell_{2}(\nu^{\star})|\leq\chi^{2}_{1,0.95}/2$. Note that $\ell_{1}(\alpha^{\star})=\ell_{2}(\nu^{\star})$ equals the minima obtained for the negative log-likelihood at the point $(\alpha^{\star},\nu^{\star})$ as well. This yields the intervals $\alpha\in[4.80,5.73]$ and $\nu\in[9.51,11.98]$ at a $95\%$ confidence level, which are consistent with their counterpart in Table 1 for $T=100$ and $n_{A}^{0}=100$.

### v2

@@ 3.2 Out-of-equilibrium probability distribution
where $p_{A}=\nu/(\nu+Q)$ and $p_{j}=\theta_{j}/(\nu+Q)$, which are precisely the probabilities given by Eqs. (39) and (40) in the particular case of equal relaxation rates. Therefore, in this limit we obtain the generating function (67) of the multinomial distribution $\mathcal{M}(n_{A}^{0},\{p_{A},p_{1},\dots,p_{S}\})$ with $n_{A}^{0}$ trials. In sum, what we have shown here is that, in the case of equal handling times, and starting by an initial population formed solely by $n_{A}^{0}$ searching consumers, the joint probability distribution of all behavioral types is multinomial in the transient as well as at stationarity. Consistently, the probabilities $p_{j}(t)$ derived in the out-of-equilibrium regime, given by Eq. (68), satisfy $p_{j}(t)\to p_{j}$ ($j=1,\dots,S$) and $p_{0}(t)\to p_{A}$ in the limit $t\to\infty$, for $p_{A}$ and $p_{j}$ given by Eqs. (39) and (40), respectively.
Out of equilibrium, feeding rates are perfectly identifiable, as can be seen from the generating function by computing an exhaustive summary. We assume that resource densities are known in advance and not subject to MLE, as well as the total number of consumers $n_{A}^{0}$. The generating function (67) is completely determined once the $S+1$ parameter combinations gathered in vector $\bm{\kappa}(\bm{\theta})=\left(\nu+\sum_{k=1}^{S}\theta_{k},\frac{\theta_{1}}{\nu},\dots,\frac{\theta_{S}}{\nu}\right)$ are known, as can be seen immediately from Eq. (68). Therefore, the exhaustive summary determines uniquely the $S$ attack rates as $\alpha_{i}=\theta_{i}/x_{R_{i}}$ ($i=1,\dots,S$) and the reciprocal mean handling time, $\nu$. Thus, the model is structurally identifiable out of equilibrium.
4 Effect of steady-state and out-of-equilibrium sampling
The two solvable models discussed above (Sections 2 and 3) illustrate the issue of parameter redundancy when steady-state probability distributions are used: in that regime, any likelihood function based on steady-state distributions will allow only for precise estimation of ratios of rates, not the rates themselves. Therefore, we have to use time-dependent probability distributions to ensure structural identifiability. However, as opposed to structural issues in model design, limitations in the dataset often preclude the whole parameter set to be inferred from data —this is known as practical identifiability (Wieland et al., 2021). In this section we focus on the effect of the quality of the data: is the problem of identification failure fully solved if we retain the temporal dependence in the probability $P_{\bm{n}}(t)$, regardless of the times at which model samples are taken?
To estimate parameters through maximum likelihood, we think of a procedure based on tracking population abundances in in silico (simulated) model samples (we name this methodology as “feeding experiment”). We want to estimate unique values for feeding rates based only on feeding experiments, in which one may have access to the population abundances of each behavioral type (in our case, the number of free consumers, $n_{A}$, and the number of consumers handling resource $i$, $n_{i}$, for $i=1,\dots,S$). We distinguish between two situations: (i) abundance data is available for different independent replicas of the system at different times at the steady state, and (ii) abundances for different behavioral types are available for out-of-equilibrium independent replicas of the system, taken at different finite times along the transient regime before reaching stationarity. We use the distributions computed in Section 3 to infer, if possible, unique values for attack rates and reciprocal mean handling times.
@@ 4 Effect of steady-state and out-of-equilibrium sampling
In both cases, our in silico feeding experiments comprise $n_{A}^{0}$ individual consumers feeding on $S$ different resources at controlled resource densities ($x_{R_{i}}$ constant, $i=1,\dots,S$), under Holling type II stochastic dynamics —cf. reaction schemes in Eqs. (9)–(10). The parameter values that we would like to estimate from abundance data are: attack rate $\alpha_{i}$ for each resource, as well as the reciprocal mean handling time over each resource, $\nu_{i}$ (relaxation rate). Assume a fixed system volume $N$. Suppose we have a detection device that can record the number of free consumers, $n_{A}(t_{j})$, as well as the number of consumers handling on each resource $i$, $n_{i}(t_{j})$ ($i=1,\dots,S$), at time $t_{j}$. In order to use the multinomial distribution (33) with probabilities (68) as the probability of any configuration of the system, every replica of the experiment starts initially with $n_{A}(0)=n_{A}^{0}$ free consumers, and no consumer is handling any prey. From now on, for the sake of simplicity, we consider that all predator individuals have equal attack and relaxation rates across resources ($\alpha:=\alpha_{1}=\dots=\alpha_{S}$, $\nu:=\nu_{1}=\dots=\nu_{S}$). Therefore, only two parameters $\alpha$ and $\nu$ are to be estimated.
We replicate the experiment a number of times $T$, each replica simulating the stochastic dynamics (using Gillespie (1977) method) with $n_{A}^{0}$ free predators initially, and record abundances at time $t_{j}$ for each realization $j=1,\dots,T$. As shown in B, it is important that at least two sampling times are different (i.e., $t_{j}\neq t_{k}$ for some $j\neq k$). Otherwise the model is not identifiable. In practice, from an empirical point of view, it is very unlikely that independent system samples are obtained at precisely the same exact time.
Thus, we have $T$ independent stochastic realizations of the system, each replica is an observation at a different time. The observational data is the following matrix
$M(\{t_{j}\})=\begin{pmatrix}n_{1,1}(t_{1})&n_{1,2}(t_{2})&\cdots&n_{1,T}(t_{T})\\ \vdots&\vdots&\ddots&\vdots\\ n_{S,1}(t_{1})&n_{S,2}(t_{2})&\cdots&n_{S,T}(t_{T})\end{pmatrix},$ (73)
where columns comprise vectors $\bm{n}_{j}=(n_{1,j},\dots,n_{S,j})$ reporting observed abundances of handling predators at sampling time $t_{j}$. Free predator abundance at time $t_{j}$ is computed with $n_{A}(t_{j})=n_{A}^{0}-\sum_{i=1}^{S}n_{i,j}(t_{j})$. If each column of $M(\{t_{j}\})$ is a sample of the Holling type II multi-resource dynamics, then the probability $P_{\bm{n}_{j}}(t_{j}|\bm{\theta})$, computed using Eqs. (33) and (68), is the probability of observing a population vector $(n_{A}(t_{j}),\bm{n}_{j})$ given a model parameter vector, $\bm{\theta}:=(\alpha,\nu)$ —notice that resource densities, $x_{R_{i}}=n_{R_{i}}/N$, the system volume, $N$, and the total number of consumers, $n_{A}^{0}$, are kept constant by assumption and therefore are not regarded as parameters in $\bm{\theta}$. Matrix $M(\{t_{j}\})$ gathers $T$ independent observations, so the function that assigns a linkelihood to parameter vector $\bm{\theta}$ given the data matrix $M(\{t_{j}\})$ is multiplicative:
$\mathcal{L}(\bm{\theta}|M(\{t_{j}\}))=\prod_{j=1}^{T}P_{\bm{n}_{j}}(t_{j}|\bm{\theta}).$ (74)
The log-likelihood function,
$\log\mathcal{L}(\bm{\theta}|M(\{t_{j}\}))=\sum_{j=1}^{T}\log P_{\bm{n}_{j}}(t_{j}|\bm{\theta}),$ (75)
@@
Figure 1: Parameter estimation using equilibrium samples. We generated a single realization of the data matrix $M(\{t_{j}\})$, including $T=100$ independent model realizations starting from $n_{A}^{0}=100$ searching predators at $t=0$, using Gillespie’s stochastic simulation algorithm up to $T=100$ different ending times uniformly drawn in the interval $[50\tau,100\tau]$ to ensure stationarity, with $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ the characteristic time of the dynamics. True rates are $\alpha=5$, $\nu=10$, and individual consumers feed on $S=5$ resources with densities $\bm{x_{R}}=(0.9,0.7,0.5,0.3,0.1)$. Left: for that synthetic dataset, we run a L-BFGS-B optimization algorithm for the log-likelihood function (75) starting from different initial conditions, and we recorded the values of $(\alpha^{\star},\nu^{\star})$ at which the maximum was reached. Optimal parameter values follow a perfect straight line. The distributions of $\alpha^{\star}$ and $\nu^{\star}$ are compatible with uniform distributions, and the optimal values of $\log\mathcal{L}$ keep constant up to round-off errors. Overall, this indicates a line of infinite optima. Right: we replicated the previous experiment 200 times and, for each replica, we estimated the ratio $\nu/\alpha$ by fitting the slope of the straight line. We obtained the distribution in the right panel, which is compatible with a Gaussian with mean 2 and variance extracted from data. The vertical, dashed line indicates the true ratio.
@@
Now, consider the effect of sampling times and assume that matrix $M(\{t_{j}\})$ is obtained when $t_{j}$ are “large enough” compared to the characteristic time of the dynamics, which is given by $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ —see Eq. (68). In Fig. 1, in silico data (i.e., abundances of each behavioral type) were generated by stochastic simulation of a system with $S$ resources (with fixed densities $x_{R_{i}}$) consumed by $n_{A}^{0}$ predator individuals. A number $T$ of independent stochastic simulations started with initial conditions $n_{A}=n_{A}^{0}$ and $n_{i}=0$ ($i=1,\dots,S$) at $t=0$, each of which were stopped up to a time $t_{\text{end}}$ enough to reach stationarity. We maximized the log-likelihood function (75) using the L-BFGS-B optimization scheme (Zhu et al., 1997), which was used for numerical optimization throughout the contribution. Each realization of the data matrix $M(\{t_{j}\})$ leads to a line of maximum likelihood in the $(\alpha,\nu)$ space (see also Fig. 2, left panel), whose slope can vary over different replicas but can be determined precisely, providing to a consistent, unbiased estimation of the ratio $\widetilde{\nu}:=\nu/\alpha$ (Fig. 1, right panel). Using data sampled at stationarity, even if ML is based on a time-dependent log-likelihood, only parameter ratios are identifiable.
@@
From Eqs. (41)–(42), it is immediate to check that the dimensionless ratio $\widetilde{\nu}$ satisfies
$\widetilde{\nu}=\frac{\langle n_{A}\rangle}{\langle n_{AR}\rangle}\sum_{i=1}^{S}x_{R_{i}},$ (76)
@@
with $\langle n_{AR}\rangle:=\left\langle\sum_{i=1}^{S}n_{i}\right\rangle=\sum_{i=1}^{S}\langle n_{i}\rangle$ the average total number of handling consumers. In B we show that the likelihood function does exhibit a line of infinite critical points with slope given by (76), in which quantities $\langle n_{A}\rangle$ and $\langle n_{AR}\rangle$ are substituted by averages across samples, $\langle n_{A}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{A,k}$ and $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}\sum_{i=1}^{S}n_{i,k}$, respectively. Therefore, our numerical experiment (Fig. 1 and Fig. (2), left panels) is consistent with a submanifold of degenerated log-likelihood optimum values in the parameter space.
@@
Figure 2: Heatmaps of negative log-likelihood. Representation of $-\log\mathcal{L}$, cf. Eq. (74), in the $(\alpha,\nu)$ space, obtained for two sets of $T=100$ model samples generated for $(\alpha,\nu)=(5,10)$: (i) taken at stationarity ($t_{\text{end}}\in[10\tau,20\tau]$, left panel), and (ii) taken out of equilibrium ($t_{\text{end}}\in[0,10\tau]$, right panel), $\tau$ being the characteristic time scale. Left: a line of infinite minima is obtained. Right: The arrow marks the true pair $(\alpha,\nu)=(5,10)$ used for data generation, and the white point marks the minimum of $-\log\mathcal{L}$. In both panels, remaining parameter values are the same as in Fig. 1.
@@
Figure 3: Distributions of parameter estimates. (a) The estimation procedure was repeated three times, for $n_{A}^{0}\in\{20,100,500\}$, using $T=20$ samples in simulated out-of-equilibrium data. Upper panels show the distributions of $\alpha^{\star}$ and $\nu^{\star}$, which tend to concentrate around true values (dashed lines) as $n_{A}^{0}$ increases. Parameter estimates show bias for low $n_{A}^{0}$ values, and the bias reduces as $n_{A}^{0}$ increase. This phenomenon is more clearly observed in the density plots represented in lower panels for $(\alpha^{\star},\nu^{\star})$, the lighter the color indicating the higher density of estimated pairs. (b) Same as (a) for $T=100$ temporal samples. Compared to panel (a), we observe that variability is decreased as $T$ increases. In both panels, remaining parameters are the same as in Fig. 1. $3000$ estimation experiments were conducted to obtain distributions.
@@
ML estimation techniques produce proper estimates for individual rates $\alpha$ and $\nu$ when data sample is taken out of equilibrium (and, importantly, at least two sampling times are different; see B). During the transient regime, samples were taken at times $\{t_{j}\}$ drawn from a uniform distribution over the interval $[0,10\tau]$, $\tau=(\nu+\alpha\sum_{i=1}^{S}x_{R_{i}})^{-1}$ being the characteristic dynamical time scale. Figure 2 compares the log-likelihood landscapes obtained using steady-state (left panel) and out-of-equilibrium model samples. A parameter screening in the $(\alpha,\nu)$ plane shows a line of infinite maxima $\nu^{\star}\propto\alpha^{\star}$ at equilibrium, confirming the results reported in Fig. 1. This evidences parameter redundancy when samples are taken at equilibrium. On the contrary, for out-of-equilibrium sampling, a single log-likelihood maximum value arises and parameters can be truly inferred from data. The accuracy of the estimates $\alpha^{\star}$ and $\nu^{\star}$ depends on data quality. For example, if the number of samples $T$ taken is large, then it is expected that estimates will be close to actual values. Similarly, if the overall number of individuals $n_{A}^{0}$ monitored in the experiment is large, then better estimates for model parameters are expected.
@@
Usually, one can only have access to one single realization of empirical observations. In this case, we still can obtain ML parameter estimates as well as confidence intervals following the log-likelihood ratio test (Cox and Snell, 1989). We illustrate this procedure using the same data matrix as in Fig. 2 (right panel), for which the ML parameter estimates are $(\alpha^{\star},\nu^{\star})=(5.24,10.65)$. We obtained log-likelihood profiles shown in Fig. 4 by minimizing negative log-likelihood over $\nu$ for fixed $\alpha$, $\ell_{1}(\alpha):=-\min_{\nu}\log\mathcal{L}(\alpha,\nu|M(\{t_{i}\}))$ (left panel), and minimizing over $\alpha$ for fixed $\nu$, $\ell_{2}(\nu):=-\min_{\alpha}\log\mathcal{L}(\alpha,\nu|M(\{t_{i}\}))$ (right panel). Then, the $95\%$ confidence interval for $\alpha$ can be defined as the set of values satisfying $|\ell_{1}(\alpha)-\ell_{1}(\alpha^{\star})|\leq\chi^{2}_{1,0.95}/2$, $\chi^{2}_{1,0.95}$ being the $95\%$ percentile of the chi-squared distribution with one degree of freedom. Similarly, the $95\%$ confidence interval for $\nu$ is defined as the set of values satisfying $|\ell_{2}(\nu)-\ell_{2}(\nu^{\star})|\leq\chi^{2}_{1,0.95}/2$. Note that $\ell_{1}(\alpha^{\star})=\ell_{2}(\nu^{\star})$ equals the minima obtained for the negative log-likelihood at the point $(\alpha^{\star},\nu^{\star})$ as well. This yields the intervals $\alpha\in[4.80,5.73]$ and $\nu\in[9.51,11.98]$ at a $95\%$ confidence level, which are consistent with their counterpart in Table 1 for $T=100$ and $n_{A}^{0}=100$.
Figure 4: Negative log-likelihood profiles for the same synthetic matrix $M(\{t_{j}\})$ used for the heatmap in Fig. 2 (right panel). Negative log-likelihood profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ (left and right panels, respectively) yield confidence intervals as the abscissae of the intercept of the profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ with the horizontal dotted lines whose ordinates are equal to $\ell_{1}(\alpha^{\star})+\chi^{2}_{1,0.95}/2$ and $\ell_{2}(\nu^{\star})+\chi^{2}_{1,0.95}/2$, respectively. $(\alpha^{\star},\nu^{\star})$ refers to the point (marked with vertical, dotted lines) at which the maximum of Eq. (74) is reached.